OpenStudy (anonymous):

How long would it take $3,100 to grow to $9,600 if the annual rate is 5.9% and interest in compounded monthly?

OpenStudy (anonymous):

solve \[9600=3100(1+\frac{.059}{12})^{12t}\] for t

OpenStudy (anonymous):

\[\frac{96}{31}=(1+\frac{.059}{12})^{12t}\] \[\frac{\ln(\frac{96}{31})}{\ln(1+\frac{.059}{12})}=12t\]

OpenStudy (anonymous):

whats the answer?